(1 point) Given the matrix [2 A = 4 (1 3 -3 5 3 1 5 , -3] find its determinant. The determinant is
Answer #1

2 4 I 3 -3 5 3 5 -3 1-3 5 Tat: 1-3 51-24 51 1+ 3 15 -3 -3 1 5 = 2(9-25) -3(-12-5) + 3 (20+3) = 2 (-16) - 3 (-17) +3 (23) -32